# A replica analysis of under-bagging

### v1

@@ 4.1 US and UB
Figure 2 shows the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and minority (positive) classes. Each panel corresponds to different values of the variance $\Delta$ and the regularization parameter $\lambda$. Recall that the $F$-measure is uppder bounded by unity. In UB, the $F$-measure increases monotonically as the number of excess data points in the majority class $\alpha^{-}-\alpha^{+}$ increases, despite of the large class imbalance. However, in the case of US, the $F$-measure does not depends on $\alpha^{-}-\alpha^{+}$, i.e., it is not affected by the class imbalance, but cannot utilize the excess data points in the majority class. This tendency does not depend on the values of $(\Delta,\lambda)$. Actually, when subsampling is used as the resampling method, one can show that the performance of the classifier does not depends on the $\alpha^{-}-\alpha^{+}$ as shown below.
@@
Let $\bar{q}=q+v$ and $\hat{\bar{\chi}}=\hat{\chi}+\hat{v}$. Then, the set of quantities $\bar{q},m,\chi$ and $\hat{Q},\hat{\bar{\chi}},\hat{m}$ are given as the solution of the following set of modified self-consistent equations. Let $\bar{\mathsf{w}}$ and $\bar{\mathsf{u}}^{\pm}$ be the solution of the following one-dimensional randomized optimization problems:
@@
$\bar{h}_{w}$ $=\hat{m}+\sqrt{\hat{\chi}}\xi_{w},\quad\xi_{w},\sim{\mathcal{N}}(0,1),$ (40)
@@
It is clear that the above modified self-consistent equation only depends on $\alpha^{+}$ and does not depends on $\alpha^{-}-\alpha^{+}$. Therefore, the $F$-measure of US cannot be improved by the excess examples the majority class111 This is exactly the same with the generalization error for the linear classifier for the data generated from symmetric Gaussian mixture model obtained in (Mignacco et al., 2020). .
@@
Figure 3: The heatmap plot for the relative log-$F$-measure $\log_{10}({\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm US})$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm US}$ are the $F$-measure for UB and US, respectively. Each panel corresponds to different values of $(\Delta,\lambda)$.
Figure 3 shows heatmap of the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm US}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm US}$ are the $F$-measure for UB and US, respectively. It is clear that the values of the $F$-measures are improved when the variance is small, the size of the minority class is small and the excess size of the majority class is large. When the variance is large, the improvement at small $\alpha^{+}$ region is small because $m$ in (36) can be small due to the large overlap between two clusters. It is also noteworthy that the improvement may be large around some values of $\alpha^{+}$ when the regularization parameter is extremely small. This indicates that the UB is robust to the existence of the phase transition from linearly separable to inseparable training data, although the performance of standard interpolator obtained by a single realization of training data is affected by such a phase transition as reported in (Mignacco et al., 2020).
@@ 4.2 UB and SW
Figure 4: The comparison of the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and the minority (positive) classes, between weighting (dashed) and UB (solid). Each panel corresponds to different values of $(\Delta,\lambda)$. Different colors of the lines represent the result for different sizes of the minority classes $\alpha^{+}$.
In this subsection, we compare the performance of UB with SW, through the values of $F$-measures. For SW, the standard weighting coefficients $\gamma^{+}=(\alpha^{+}+\alpha^{-})/(2\alpha^{+}),\gamma^{-}=(\alpha^{+}+\alpha^{-})/(2\alpha^{-})$ are used.
Figure 4 shows the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and minority (positive) classes. In contrast to US and UB, the $F$-measure of the SW method degrades as the number of the excess examples of the majority class increases. In other words, as the amount of the class imbalance increases, the prediction performance of the estimator obtained by simple class-wise weighted cost degrades, although the degradation is mild if the size of the minority class is large. Similarly, Figure 4 shows heatmap of the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm SW}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm SW}$ are the $F$-measure for UB and SW, respectively. When the size of the minority class is sufficiently large, the performance is almost comparable. However, when the size of the minority class is small and the class imbalance is large, the performance of the simple weighting method is several times to tens of thousands of times worser than UB in terms of the value of $F$-measure.
@@
Figure 5: The heatmap plot for the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm weighting}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm weighting}$ are the $F$-measure for UB and the weighting method, respectively. Each panel corresponds to different values of $(\Delta,\lambda)$.
@@ 5 Summary and Discussions
By using the derived sharp asymptotics, we investigated the performance of the UB, US and SW. Our main findings were (i) UB can improve the performance of the classifier in terms of the $F$-measure even if the class-imbalance is huge, (ii) the performance of the classifier obtained by US does not depends on the size of the excess majority examples, and (iii) the performance of the SW method degrades as the size of the excess majority examples increases especially when the size of the minority class is small and the class imbalance is large. Furthermore, UB seemed to be robust to the interpolation phase transition in contrast to US.
Although UB is a powerful method for learning in imbalanced data, but it requires an increased computational cost proportional to the number of under-sampled datasets, which can be a serious issue in learning deep neural networks. Therefore, developing simple heuristics to achieve similar results as UB would be a promising future direction.

### v2

@@ 4.1 US and UB
Figure 3 shows the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the scaled size between the majority (negative) and minority (positive) classes, or equivalently, the scaled size of the excess majority examples. Each panel corresponds to different values of the variance $\Delta$ and the regularization parameter $\lambda$. Recall that the $F$-measure is uppder bounded by unity. In UB, the $F$-measure increases monotonically as the number of excess data points in the majority class $\alpha^{-}-\alpha^{+}$ increases, despite of the large class imbalance. However, in the case of US, the $F$-measure does not depends on $\alpha^{-}-\alpha^{+}$, i.e., it is not affected by the class imbalance, but cannot utilize the excess data points in the majority class. This tendency does not depend on the values of $(\Delta,\lambda)$. Actually, when subsampling is used as the resampling method, one can show that the performance of the classifier does not depends on $\alpha^{-}-\alpha^{+}$ as shown below.
@@
Let $\bar{q}=q+v$ and $\hat{\bar{\chi}}=\hat{\chi}+\hat{v}$. Then, the set of quantities $\bar{q},m,\chi$ and $\hat{Q},\hat{\bar{\chi}},\hat{m}$ are given as the solution of the following set of modified self-consistent equations when subsampling (6)-(7) is used for the resampling method. Let $\bar{\mathsf{w}}$ and $\bar{\mathsf{u}}^{\pm}$ be the solution of the following one-dimensional randomized optimization problems:
@@
$\bar{h}_{w}$ $=\hat{m}+\sqrt{\hat{\bar{\chi}}}\xi_{w},\quad\xi_{w},\sim{\mathcal{N}}(0,1),$ (40)
@@
It is clear that the above modified self-consistent equation only depends on $\alpha^{+}$ and does not depends on $\alpha^{-}-\alpha^{+}$. Therefore, the $F$-measure of US cannot be improved by the excess examples the majority class222 Actually, this is exactly the same with the equations to determine the behavior of linear classifiers trained on the data generated from symmetric Gaussian mixture model with a data size $\alpha^{+}N$ and an input dimension $N$ obtained in (Mignacco et al., 2020). .
@@
Figure 4: Heatmap plot for the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm US}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm US}$ are the $F$-measure for UB and US, respectively. Each panel corresponds to different values of $(\Delta,\lambda)$.
Figure 4 shows heatmap of the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm US}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm US}$ are the $F$-measure for UB and US, respectively. It is clear that the values of the $F$-measures are improved when the variance is small, the size of the minority class is small and the excess size of the majority class is large. When the variance is large, the improvement at extremely small $\alpha^{+}$ region is small because $m$ in (36) can be small due to the large overlap between two clusters.
4.1.1 Double-descent-like behavior
(a) $(\sqrt{\Delta},\lambda)=(1.5,10^{-5})$
(b) $(\sqrt{\Delta},\lambda)=(0.75,10^{-5})$
(c) $(\sqrt{\Delta},\lambda)=(0.5,10^{-5})$
(d) $(\sqrt{\Delta},\lambda)=(1.5,10^{-5})$
(e) $(\sqrt{\Delta},\lambda)=(0.75,10^{-5})$
(f) $(\sqrt{\Delta},\lambda)=(0.5,10^{-5})$
Figure 5: Relative variance $v/(q+v)$. Upper panels: the relative variance shown as a function of $\alpha^{+}$ for some selected values of $\alpha^{-}-\alpha^{+}$. The legend is common across all panels. Red dashed lines represent the interpolation threshold. Lower panels: the relative variance shown as a 2d-heatmap.
It is also noteworthy that the improvement may be large around some values of $\alpha^{+}$ when the regularization parameter is extremely small. This is because the variance associated with random reweighting coefficient $\mathbf{\boldsymbol{c}}$ exhibits double-descent-like behavior with respect to $\alpha^{+}$ and takes a large value at a certain $\alpha^{+}$. This point is evident in Figure 5 that shows the relative variance $v/(q+v)$ for several values of $\Delta$. It is clear that the relative variance takes a peak value at the interpolation threshold obtained by the formula in the logistic regression with balanced two-component Gaussian mixture with data size $\alpha^{+}N$ (Mignacco et al., 2020). Similar double-descent-like behavior is also observed in an analysis of bagging without class imbalance structure (Clarté et al., 2024). This indicates that the UB, which can remove the contribution from $v$, is robust to the existence of the phase transition from linearly separable to inseparable training data, although the performance of standard interpolator obtained by a single realization of training data is affected by such a phase transition as reported in (Mignacco et al., 2020).
@@ 4.2 UB and SW
Figure 6: The comparison of the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and the minority (positive) classes, between weighting (dashed) and UB (solid). Each panel corresponds to different values of $(\Delta,\lambda)$, where $\lambda$ is the regularization parameter for UB. Different colors of the lines represent the result for different sizes of the minority classes $\alpha^{+}$.
In this subsection, we compare the performance of UB with SW, through the values of $F$-measures. For SW, the standard weighting coefficients $\gamma^{+}=M/(2M^{+})=(\alpha^{+}+\alpha^{-})/(2\alpha^{+}),\gamma^{-}=M/(2M^{-})=(\alpha^{+}+\alpha^{-})/(2\alpha^{-})$ are used.
Figure 6 shows the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and minority (positive) classes. In contrast to US and UB, the $F$-measure of the SW method decreases as the number of the excess examples of the majority class increases. In other words, as the amount of the class imbalance increases, the prediction performance of the estimator obtained by simple class-wise weighted cost deteriorates, although the amount of the decrease is mild if the size of the minority class is large. Similarly, Figure 6 shows heatmap of the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm SW}$ in a log-scale, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm SW}$ are the $F$-measure for UB and SW, respectively. When the size of the minority class is sufficiently large, the performance is almost comparable. However, when the size of the minority class is small and the class imbalance is large, the performance of the simple weighting method is several times to tens of thousands of times worser than UB in terms of the value of $F$-measure.
@@
Figure 7: The heatmap plot for the relative $F$-measure in log-scale: $\log_{10}{\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm weighting}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm weighting}$ are the $F$-measure for UB and the weighting method, respectively. Each panel corresponds to different values of $(\Delta,\lambda)$, where $\lambda$ is the regularization parameter used in UB.
@@ 5 Summary and Discussions
Using the derived sharp asymptotics, we investigated the performance of the UB, US, and SW. Our main findings were (i) UB can improve the performance of the classifier in terms of the $F$-measure by increasing the size of the majority class while keeping the size of the minority class fixed, even when the degree of class imbalance becomes large, (ii) the performance of the classifier obtained by US does not depend on the size of the excess majority examples, and (iii) the performance of the SW method deteriorates as the size of the excess majority examples increases especially when the size of the minority class is small and the class imbalance is large. Furthermore, UB was robust to the interpolation phase transition in contrast to US.
In conclusion, the ensembling considering the class imbalance structure is different from the combination of US and ridge regularization. This result is clearly different from the case of naive bagging in training GLMs without considering the class imbalance structure, where ensembling and ridge regularization give the same performance.
Although UB is a powerful method for learning in imbalanced data, it requires an increased computational cost proportional to the number of under-sampled datasets, which can be a serious problem in learning deep neural networks. Therefore, the development of simple heuristics to achieve similar results as UB would be a promising future direction.